The earth exerts a gravitational force of 1,9 x 1020 N on the moon. The mass of the moon is 7,4 x 1022 kg and its radius is 1,74 x 106 m.



Calculate the magnitude of the gravitational acceleration on the moon

1
Expert's answer
2021-11-25T16:22:15-0500

"a = \frac{F}{m_m} \\

a = \frac{1.9 \times 10^{20}}{7.4 \times 10^{22}} = 0.256 \times 10^{-2} \\

a = 2.56 \times 10^{-3} \; m/s^2"